OpenStudy (anonymous):

what point does this line pass through, which is the basis of this equation y-3=3(x+1) is the answer (-1,3)

OpenStudy (shamil98):

yep, you got it right.

OpenStudy (shamil98):

y -y1 = m(x-x1) y - 3 = 3(x+1) (-1,3) is a point on the line in which it passes through

OpenStudy (anonymous):

thank you I was unsure